OpenStudy (anonymous):

A sample of a volatile liquid is vaporized completely in a 240.0 cm3 flask at 99°C and 757 mm Hg. The condensed vapor weighs 0.465 g. What is the molar mass of the liquid?

OpenStudy (zpupster):

PV=nRt or n=PV/RT n= no. of moles now convert the units to match R's unit R the ideal Law gas constant = 8.31 J/KT so remember J=m^3*Pa Pressure=(in Pa) 133.2*757mmHg = Volume =(in m^3) = 240*10^-6 m^3= R=8.31 J/KT T (in K) = 99C+273 = place this in our formula and we will have our moles = ?moles molar mass = 0.465g./?moles

OpenStudy (anonymous):

so it's the same as the mass of the condensed vapor?

OpenStudy (joannablackwelder):

Yes. Mass is conserved.

OpenStudy (anonymous):

oh thanks you

OpenStudy (joannablackwelder):

You're welcome.

OpenStudy (anonymous):

oh wait i was only able to do conversion at operation as of now. and i got 0.0078. hmmmm why??

OpenStudy (joannablackwelder):

Did you be careful to use consistent units with your R value?

OpenStudy (anonymous):

yes i actually tried it twice. same outcome though

OpenStudy (joannablackwelder):

I converted to Liters, Kelvin, and atm and used 0.0820574 for R. I got n=0.00783 moles. Then mass/moles gives molar mass.

OpenStudy (anonymous):

ohhh exactly the same haha. mass of what should be used? in mass/moles :)

OpenStudy (joannablackwelder):

:) the mass given in the problem statement.

OpenStudy (anonymous):

oh right my bad haha :) so 59.62 g/mol

OpenStudy (joannablackwelder):

Yes, that is what I get. :)

OpenStudy (anonymous):

wahha thankyou :)))

OpenStudy (joannablackwelder):

Good job. :)

OpenStudy (joannablackwelder):

No worries.